OpenStudy (anonymous):

solve the system of linear equations by graphing. 2x+y=0 and 5x+y=3

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

It's easy to graph this equations when we put them in slope intercept form,and that is, \[\large \color{blue}{y=mx+b}\] Let's start with the first equation \[\large 2x+y =0\] subtract 2x from both sides. \[\large -2x + 2x +y = -2x+0 \]Simplify \[\large y= -2x + \color{blue}0 \] now it matches out slope intercept form equation in that the "slope" is \[\large m=-2\]and our b= 0. Now the second equation. \[\large 5x+y = 3\] We can put this into slope intercept form as well. subtract 5x from both sides \[\large -5x + 5x +y = -5x + 3\]Simplify \[\large y = -5x + \color{blue}3\] In this equation,our "b" = 3 and our slope is -5

OpenStudy (anonymous):

ok so would the ordered pair then be (3,-5)?

OpenStudy (anonymous):

the two order pairs are (-2,0) and (-5,3) am I correct?

OpenStudy (amistre64):

your x should equal 1 .... and then you can find y the point is that a solution to a system of equation can be approximated by finding the point(s) of intersection on a picture

OpenStudy (jhannybean):

I'm guessing you're looking for their point of intersection. Yes?

OpenStudy (jhannybean):

Ah, ok.

OpenStudy (anonymous):

I have to solve this system of linear equations, then write what the ordered pair is and then graph it on a graph

OpenStudy (amistre64):

oh, becasue you posted: "solve the system of linear equations by graphing."

OpenStudy (anonymous):

sorry that is what it says... but then under it says use the graphing tool to graph the system. What is the solution of the system of equations? Select the correct choice below and if necessary fill in the answer box to complete your choice.

OpenStudy (amistre64):

you are going to have to use the graphing tool to plot 2 points and line them in 

OpenStudy (anonymous):

the choices are A. ___________(type an ordered pair) B. there are infinitely many solutions C. there is no soluton

OpenStudy (jhannybean):

use the two lines we solved for and graph them,then find the intersecting point

OpenStudy (anonymous):

but how would i graph them